A cannon of mass M (=1300kg) fires a 72-kg ball in a horizontal direction with speed v (= 55 m/s) relative to the cannon. The cannon is mounted so that it can recoil freely.
What is the velocity V of the recoiling cannon with respect to Earth?
What is the initial velocity vE of the ball with respect to Earth?. 
 

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Dear Student

Assuming that the projectile speed is measured from a fixed point equal to the tip of the muzzle of the cannon (and not the cannon itself)


First you would find the momentum of the cannon ball and since both the cannon and the ball would have equal but opposite momentums you can then solve -

72kg * 55 m/s = 1300kg * X m/s

3960kgm/s = 1300kg * X m/s

3960kgm/s / 1300kg = X m/s

3.046 m/s = X m/s


Thanks